OpenStudy (anonymous):

a boy throws stone with a speed of 12m/s if he releases them at a height 1.25m the maximum distance from his feet where they land is?

OpenStudy (anonymous):

i got the answer as 14.4m but the orginal answer is 15.6m...

OpenStudy (anonymous):

i used Rmax=u^2/g

OpenStudy (anonymous):

So he threw the stone vertically up?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

If I use the equation \[d = \frac{v^2}{g}\sin(2\theta)\] and set theta to 45, then I get 14.4

OpenStudy (anonymous):

i think we have to add 1.25m also

OpenStudy (anonymous):

right, but that gives 15.65 instead of just 15.6

OpenStudy (anonymous):

can we make a approximate answer

OpenStudy (anonymous):

I don't think so. I think the 15.6 answer came from some working out :)

OpenStudy (anonymous):

I used \[y = v_0\sin(\theta)t - \frac{1}{2}gt^2\]and solved for t with y = -1.25, v_0 = 12, g = 10, and theta = 45 degrees. Then I took the positive solution and plugged it into\[ x = v_0\cos(\theta)t\]to get something like 15.55 m

OpenStudy (anonymous):

something like \[x = \frac{3(\sqrt{194} + 12)}{5}\text{m}\]

OpenStudy (anonymous):

let me solve this and i will inform u.......

OpenStudy (anonymous):

wat is the t?

OpenStudy (anonymous):

After solving that first equation, I got a negative t, and\[t = \frac{\sqrt{97} + 6\sqrt{2}}{10}\]